At a pressure of 405 kPa, the volume of a gas is 6.00 cm3. Assuming the
temperature remains constant, at what pressure will the new volume be 4.00
cm3?
1
Expert's answer
2013-05-02T11:18:16-0400
pV = nRT
p - pressure, V - volume, n - the amount of moles, R - gs constant, T - temperature.
The gas is unknown.
In this task we speak about the same amount of gas but under different pressure. We can assume that it's mass does not change. The amount of moles is proportional to mass. That's why the amount of moles is constant too. So the amount of moles is constant. We can assume that the right part of the equation is constant. We can write:
p1V1 = p2V2
405*6 = x*4
x = (405*6)/4 = 607 kPa
That is why at the pressure 607 kPa the new volume will be 4.00 cm3.